### v1

# Privacy for Quantum Annealing. Attack on Spin Reversal Transformations in the case of cryptanalysis

Abstract
This paper demonstrates that applying spin reversal transformations (SRT), commonly known as a sufficient method for privacy enhancing in problems solved using quantum annealing, does not guarantee privacy for all possible problems. We show how to recover the original problem from the Ising problem obtained using SRT when the resulting problem in Ising form represents the algebraic attack on the $E_{0}$ stream cipher. A small example is used to illustrate how to retrieve the original problem from the one transformed by SRT. Moreover, it is shown that our method is efficient even for full-scale problems.
keywords:
Privacy enhancement, quantum cloud computing, Ising problem, quantum annealing, secure cloud computing
volume: 185issue: 1
Privacy for Quantum Annealing
1 Introduction
Quantum annealing is a highly complex process. In this paper, we focus only on its applications and the confidentiality of the data used in the calculations. In the following section, the aspects of the communication scheme assumed, and the calculations used are discussed.
1.1 Quantum cloud services and privacy.
Quantum cloud services are now becoming more and more popular. The delivered products allow us to solve a wide range of problems. In this paper, we focus on two specific quantum services:
  • Quantum Annealing, introduced by Tadashi Kadowaki and Hidetoshi Nishimori in [1];
  • Quantum Approximate Optimization Algorithm, introduced by Edward Farhi et al. in[2].
As presented in [3], quantum annealing has many applications. It can be used in areas such as:
  • traffic flow optimization;
  • logistic, vehicle routing problem;
  • in finance, portfolio optimization;
  • quantum simulation.
However, typically, solutions are implemented using remote access. Such a solution requires the transfer of data to an outsourcer. Under ideal conditions, such a scheme can be depicted as in the figure 1.
Figure 1: Scheme of ideal communication.
The tasks to be solved in both cases require representation using the Ising model. In this case, communication with the quantum cloud requires sending a connection matrix $J$ and a vector of biases $h$, presented as equations successively (3) and (2). The communication scheme is straightforward, as presented in the figure 1. Client prepares problem in Ising model (in the case of Quantum Annealing or Quantum Approximate Optimization Algorithm) and then sends it to Quantum Cloud Service via the internet. After computations are done, the solution is sent to the client similarly.
In specific cases, such an arrangement is unacceptable. As presented in [4], [5] or [6] quantum annealing can be used in cryptoanalysis. In such use, it is not permissible to transfer the data to an external party. The answer is privacy-preserving methods.
Privacy in quantum computing is a promising area of research. Secure protection algorithms will expand the market for quantum services. As it said in [7], Secure Quantum Computing can be divided into two groups of methods:
  • Blind Quantum Computing is a technique that allows outsourcing computations without disclosing the computation’s details to the server. This group of protocols is often impractical and requires the customer to have the quantum hardware or the participation of several servers.
  • Quantum Homomorphic is encryption that involves performing calculations on encrypted data. The result of the calculation, after decryption, is the result of the original problem.
This paper focuses on a method called spin reversal transformation (SRT) presented in [8]. The authors propose homomorphic encryption for quantum annealing, which they believe protects the details of quantum annealing instances against a malicious cloud. As said in the paper, the scheme runs on the Ising model, so it can also be used for privacy preservation in the Quantum Approximate Optimization Algorithm.
1.2 Man in the middle. The real threat.
The communication diagram shown in Figure 1 has little to do with actual communication conditions. With the growing number of cloud service providers, the threat is growing. Providers services are highly needed, and their availability is severely limited. This leads to a situation where a new provider offering competitive services can gain popularity rapidly. The internal structure of cloud services remains a mystery, and it is not known if there is no eavesdropping between the client and the actual computer. Such a new, untrusted provider may be malicious, and communications between the user and the actual quantum computer may be intercepted.
Figure 2 shows a more realistic scheme. Between client and quantum cloud service is the adversary. All communication goes through the adversary, which has any possibility to perform operations on the received problem. This paper assumes that adversaries eavesdrop on intercepted communications and pass them forward. Intercepted data can lead to financial losses (for commercial applications) or even threaten security (in the case of military applications such as cryptanalysis).
Figure 2: Communication scheme with a potential adversary.
1.3 A description of the models used
As mentioned earlier, using the Quantum Approximate Optimization Algorithm or Quantum Annealing requires presenting the problem in an sepcific form. One such form is the Ising model, known since 1920. Ernst Ising and Wilhelm Lenz introduced it as a description of magnetic materials. Despite its original application, the model has also been applied in combinatorial optimization. As presented in [9] there exist Ising formulations of many NP problems such as:
  • Graph partitioning [10];
  • Knapsack problem [11];
  • Graph coloring [11];
  • Traveling salesman [11].
Moreover, the model can be applied in cryptanalysis. Including but not limited to the factorization presented by Jiang, in [4], breaking the discrete logarithm [12, 13, 14, 15, 16] or attacking symmetric ciphers [5, 17, 6]. It is worth noting that the last two applications require a transition between the QUBO model and the Ising model.
As mentioned in [18], each Ising problem can be viewed as a minimizing expression presented as Equation (1).
$f(s)=\sum_{i=0}^{n-1}h_{i}s_{i}+\sum_{i,j=0}^{n-1}J_{i,j}s_{i}s_{j}$ (1)
Vector $s$ is called state and each variable $s_{i}\in\{-1,1\}$ is called spin. In practice, the following are used to characterize the problem:
  • a vector of biases:
    $h=\begin{bmatrix}h_{0}&h_{1}&\cdots&h_{n-1}\end{bmatrix}^{T};$ (2)
  • a matrix describing connections between variables:
    $J=\begin{bmatrix}J_{0,0}&J_{0,1}&\cdots&J_{0,n-1}\\ J_{1,0}&J_{1,1}&\cdots&J_{1,n-1}\\ \vdots&\vdots&\ddots&\vdots\\ J_{n-1,0}&J_{n-1,1}&\cdots&J_{n-1,n-1}\end{bmatrix}.$ (3)
An alternative model is the QUBO model, which uses binary variables. Quadratic unconstrained binary optimization can also be viewed as minimizing specific expression, shown as Equation (4).
$f(x)=\sum_{i=0}^{n-1}Q_{i,i}x_{i}+\sum_{i,j=0}^{n}Q_{i,j}x_{i}x_{j}.$ (4)
Vector $x$ contains $n$ binary variables.
As mentioned earlier, there is a simple transformation between the Ising model and the QUBO problem. Its main idea is shown in the Figure 3. The transition between models requires simple substitutions:
  • from Ising model to QUBO:
    $x_{i}=\frac{1}{2}\cdot(s_{i}+1),$ (5)
  • from QUBO to Ising model:
    $s_{i}=2x_{i}-1.$ (6)
In the rest of the paper, we skip transitions between models, considering the operation trivial.
Figure 3: Transformation between models.
As described in [7], coefficients of each matrix can be determined using matrix coefficients for the corresponding problem. When moving from the Ising model to the QUBO problem, the connection matrix is determined as follows:
$\begin{array}[]{c}Q_{i,j}=4J_{i,j},\\ Q_{i,i}=2(h_{i}-\sum_{j}J_{i,j}-\sum_{j}J_{j,i}.)\end{array}$ (7)
At the transition in the opposite direction, the bias vector is determined as follows:
$h_{i}=\frac{Q_{i,i}}{2}+\frac{\sum_{j}Q_{i,j}+\sum_{j}Q_{j,i}}{4}.$ (8)
The connection matrix is determined as follows:
$J_{i,j}=\frac{Q_{i,j}}{4}.$ (9)
2 Spin Reversal Transformation in detail
The simplest way to protect privacy is to encrypt the problem using a homomorphic algorithm. The first ideas of privacy protection in this way were described in [19]. However, they were not very popular for a long time. The idea of applying such a method to quantum cloud computing is described in [8] and partly in [7]. The method utilizes a random sequence to reverse a given sign in an Ising problem instance. It is important to note that the minimal energy of the instance does not change. As depicted in [8] and presented using the Equation (12), the solution to the concealed problem, when uncovered, is the solution to the original problem. In addition, the authors of the paper demonstrate that their method is efficient.
2.1 Secure communication scheme
The idea of a secure scheme is presented in Figure 4. The communication channel is secured by encrypting the problem being sent. As presented in [8], a suitable algorithm will recover the solution to the original problem from the encrypted solution. After preparing the problem, the client will perform encryption and send the new problem to the Quantum Cloud Service. An adversary, having intercepted such a problem, according to the authors, will not be able to reconstruct the original problem without knowing the original key. After processing, the solution to the encrypted problem is sent to the client. Using a previously used key, the customer can recover the solution to the original problem.
Figure 4: A secure communication scheme, proposed in [8].
2.2 Description of the algorithm
The encryption scheme described in [8] is based on SRT, called gauge transformation [20]. The transformation uses a binary string $x$ to change the signs of the selected coefficients. After applying mentioned transformation, Equation (1) is transformed as follows:
$f^{*}(s^{*})=\sum_{i=0}^{n-1}(-1)^{x_{i}}h_{i}s^{*}_{i}+\sum_{i,j=0}^{n-1}(-1)^{x_{i}+x_{j}}J_{i,j}s^{*}_{i}s^{*}_{j}.$ (10)
The same sequence $x$ must be used to determine the original solution. The corresponding solution can be determined according to the Equation 11.
$s_{i}=(-1)^{x_{i}}s_{i}^{*}$ (11)
As mentioned in [8] and [20], this transformation does not change the ground state. Furthermore, a solution $s$ minimizes $f(s)$ if and only if solution $s^{*}$ minimizes $f^{*}(s^{*})$. This follows from the equality shown in the Equation below (12)
$\begin{array}[]{c}f^{*}(s^{*})=\sum_{i=0}^{n-1}(-1)^{x_{i}}h_{i}\cdot(-1)^{x_{i}}s_{i}+\sum_{i,j=0}^{n-1}(-1)^{x_{i}+x_{j}}J_{i,j}\cdot(-1)^{x_{i}}s_{i}\cdot(-1)^{x_{j}}s_{j}=\\ \sum_{i=0}^{n-1}(-1)^{2x_{i}}h_{i}s_{i}+\sum_{i,j=0}^{n-1}(-1)^{2x_{i}+2x_{j}}J_{i,j}s_{i}s_{j}=\\ \sum_{i=0}^{n-1}h_{i}s_{i}+\sum_{i,j=0}^{n-1}J_{i,j}s_{i}s_{j}=f(s)\\ \end{array}$ (12)
The described transformation can be outlined as a simple scheme, as in [8]:
  1. 1.
    The client generates a secret key $x=(x_{0},x_{1},\ldots,x_{n-1})$;
  2. 2.
    The client computes $h^{*}$ and matrix $J^{*}$:
    $h_{i}^{*}=(-1)^{x_{i}}h_{i},$ (13)
    $J_{i,j}^{*}=(-1)^{x_{i}+x_{j}}J_{i,j},$ (14)
  3. 3.
    The client sends the concealed problem to the quantum cloud service and receives the solution $s^{*}$;
  4. 4.
    The client retrieves the solution using equation (11).
3 Details of the proposed attack
This paper argues that the privacy-enhancing mechanism presented in [8] fails in many instances. As a counterexample, we demonstrate that the method fails when the method of transforming the stream cipher cryptanalytic problem is known. Specifically, we show that, without prior knowledge of the random sequence used to conceal the given Ising problem for cryptanalysis of the $E_{0}$ cipher, one can easily retrieve the original problem using the intercepted data. Furthermore, by knowing the solution to the concealed problem, one can also deduce the solution to the original problem. It also allows the recovery of the concealment key and the data on which the client cares about confidentiality (key hidden in the solution to the problem sent, ciphertexts, or keystream). With a concealment key, the adversary can recover communications encrypted with this key.
3.1 The problem for the $E_{0}$ cipher
We will present an attack on the mentioned scheme using the QUBO problem corresponding to the cryptanalysis of the $E_{0}$ cipher, presented in the [21]. As the authors point out, the reduction presented in the paper in the next few years can be practically embedded in a commercially available quantum annealer. This opens up the possibility of application to real cryptanalysis, for which it is essential to maintain the privacy of transmitted data and recovered keys.
The cipher is built from three elements:
  • 4 shift registers with linear feedback, specified by the following primitive polynomials $f_{i}(x)$:
    $\begin{array}[]{cc}L_{1}:&f_{1}(x)=x^{25}\oplus x^{20}\oplus x^{12}\oplus x^{8}\oplus 1,\\ L_{2}:&f_{2}(x)=x^{31}\oplus x^{24}\oplus x^{16}\oplus x^{12}\oplus 1,\\ L_{3}:&f_{3}(x)=x^{33}\oplus x^{28}\oplus x^{24}\oplus x^{4}\oplus 1,\\ L_{4}:&f_{4}(x)=x^{39}\oplus x^{36}\oplus x^{28}\oplus x^{4}\oplus 1.\end{array}$ (15)
  • Summation Combiner Logic, computing the two-bit value $s_{t+1}$:
    $F_{1}:s_{t+1}=\Bigl{\lfloor}\frac{\sum_{i=1}^{4}x_{i}+c_{t}}{2}\Bigr{\rfloor}$
  • Blend Register, calculating the two-bit value of $c_{t+1}$, using the bijections described in the cipher specification:
    $F_{2}:c_{t+1}=s_{t+1}\oplus T_{1}[c_{t}]\oplus T_{2}[c_{t-1}].$
  • Each keystream bit is determined by the value from shift registers with linear feedback, $x_{0},x_{1},x_{2},x_{3}$, and one bit from the Blend Register:
    $z_{i}=x_{1}\oplus x_{2}\oplus x_{3}\oplus x_{4}\oplus c_{t}^{0}.$
Each bit of the key stream is described by 8 equations:
$\begin{cases}f_{0}:z_{t}=l_{t+1}\oplus m_{t+7}\oplus n_{t+1}\oplus o_{t+7}\oplus c^{0}_{t},\\ f_{1}:l_{t+25}=l_{t}\oplus l_{t+5}\oplus l_{t+13}\oplus l_{t+17},\\ f_{2}:m_{t+31}=m_{t}\oplus m_{t+7}\oplus m_{t+15}\oplus m_{t+19},\\ f_{3}:n_{t+33}=n_{t}\oplus n_{t+5}\oplus n_{t+9}\oplus n_{t+29},\\ f_{4}:o_{t+39}=o_{t}\oplus o_{t+3}\oplus o_{t+11}\oplus o_{t+35},\\ f_{5}:c^{1}_{t+1}=s^{1}_{t+1}\oplus c^{1}_{t}\oplus c^{0}_{t-1},\\ f_{6}:c^{0}_{t+1}=s^{0}_{t+1}\oplus c^{0}_{t}\oplus c^{1}_{t-1}\oplus c^{0}_{t-1},\\ f_{7}:4s^{1}_{t+1}+2s^{0}_{t+1}+\beta=l_{t+1}+m_{t+7}+n_{t+1}+o_{t+7}+2c^{1}_{t}+c^{0}_{t}.\end{cases}$ (16)
3.2 How to identify the problem and parameterize it?
Quantum service providers require many data to identify the user. Among such data, we can distinguish:
  • first and last name;
  • email address;
  • company name, job title, field of study;
  • the purpose of computer access.
With their use, an untrusted provider can identify the client and the research area in which the client works. Identifying the area of study will allow the data obtained to be matched with one of the known problems. Problems vary significantly in size, density, and coefficient range. Betting that resources are limited and methods cannot change the connectivity (density) or size of the problem.
The critical element of the attack is determining a parameterized matrix for the problems under consideration. It should be noted that such a matrix is determined only once for a given problem. Determination of the matrix acts as a precomputation. The interactive attack phase uses pre-made parameterized matrices.
The details of the parameterization will depend on the type of cipher. The general principle remains the same. The construction for the $E_{0}$ cipher is discussed below. As shown in Equation (16), only one of the equations depends on the bits of the keystream. According to a method of transformation algebraic attack presented in [5], for each keystream bit, the component of the final polynomial will be computed as below:
$z_{t}+l_{t+1}+m_{t+7}+n_{t+1}+o_{t+7}+c^{0}_{t}-2K=0\Big{/}^{2}$ (17)
Finally, six variables depending on keystream bits are obtained: $2z_{t}\cdot l_{t+1},2z_{t}\cdot m_{t+7},2z_{t}\cdot n_{t+1},2z_{t}\cdot o_{t+7},2z_{t}\cdot c_{t}^{0},-4z_{t}\cdot K$, note that K, is an integer variable and should be replaced with a binary variable. The coefficients resulting from the squaring of other equations are constant. According to Equations (8), equations dependent on keystream occurring in the vector of biases can be computed.
The alternative method does not require the direct analysis of equations. As shown above, each of the coefficients depends on the keystream bits linearly. A similar relationship will be satisfied for other stream ciphers.
Assume that an oracle $\phi$ is given which, for any keystream, will return an Ising model, as Equation (19). The oracle can be constructed based on known publications on selected attacks. A detailed description of the transformation of the algebraic attack to the QUBO problem, as in [5], [6], [21], allows the construction of the assumed oracle. In known cases, the construction of such an oracle does not require a large number of resources. As in the method requiring analysis of each equation, the oracle must be queried only in preparing parameterized matrices for attacks.
The parameterized matrix can be determined with such knowledge, as discussed below. Keystream with a hamming weight of 1 will be used to determine the matrix. Such stream can be demoted as $s_{i}$ and is presented as Equation (18).
[…]
We denote $h_{i}$, $J_{i}$ as Ising model, such that:
$h_{i},J_{i}\leftarrow\phi(s_{i}).$ (19)
Additionally, we denoted $h_{\infty}$, $J_{\infty}$ as the Ising model for a stream with Hamming weight equals zero. Then, the parametrized problem is computed as follows:
$\begin{array}[]{c}h_{P}=h_{\infty}+\sum_{i=0}^{n}(h_{i}-h_{\infty})\cdot z_{i},\\ J_{P}=J_{\infty}+\sum_{i=0}^{n}(J_{i}-J_{\infty})\cdot z_{i},\end{array}$ (20)
where, $z_{i}$ is variable.
3.3 Executing the attack, the most straightforward phase
Using a parameterized problem, an attack is performed. A vector of biases is sufficient to perform the attack. The resulting encrypted vector of biases is juxtaposed with a parameterized one to form a system of equations. As can be seen from the construction of the parameterized matrix, each of the equations of the system can be presented as Equation (21). In general, the coefficients of a vector can be the sum of a rational number and a linear combination of any keystream bits, as $\mathcal{A}_{i}$, we denote the set of indices of the relevant bits.
$h_{i}^{*}=\sum_{j\in\mathcal{A}_{i}}+a_{i},$ (21)
where $a_{i}\in\mathbb{Q}$ The next step is to analyze each of the resulting equations in the correct order. First, the equations must be transformed to form presented as Equation (22)
$z_{j}+b_{i}=h_{i}^{*},$ (22)
where, $b_{i}\in\mathbb{Q}$. If the transformation of the equations is not possible, reduce the system to upper triangular form. For example, using the Gaussian reduction. Then, an equation analysis is performed from the last equation, considering previous solutions. The variable $z_{j}$ is a binary variable, $z_{j}\in\{0,1\}$, therefore the relation (23) should be satisfied.
$h_{i}^{*}-b_{i}\in\{0,1\}.$ (23)
If the above relation is satisfied, then:
  • the relevant coefficient has not been concealed;
  • designated bit of concealment key is: $x_{i}=0$;
  • designated bit of keystream is: $z_{j}=h_{i}^{*}-b_{i}$.
Otherwise, the given coefficient is concealed and:
  • designated bit of concealment key is: $x_{i}=1$;
  • designated bit of keystream is: $z_{j}=-h_{i}^{*}-b_{i}$.
In most cases, solving the equation above gives the bit of the keystream $z_{j}$ with a probability of $1$. If one finds that $z_{j}\in\{0,1\}$ regardless of whether a coefficient is concealed, one can check if other equations where $z_{j}$ appears are correct. The above procedure allows the recovery of the used keystream. The remaining bits of the concealment key can be determined by comparing the remaining coefficients. If the coefficients are opposite, the given coefficient has been concealed, and the corresponding concealment key bit is 1.
For every bit of the keystream, retrieving the correct value of the bit requires only solving an affine equation, and the whole attack is swift. In addition, the attack allows the recovery of the concealment key. With the use of a key:
  • any further problem encrypted with the same key can be exposed;
  • the solution to the problem sent may be unveiled, including knowledge of the secret key, which the client wanted to recover.
4 A practical example of the proposed attack
To illustrate the correctness of the attack, an attack performed in practice is presented. A scaled-down $E_{0}$ cipher was used to show the step-by-step operation of the attack. In addition, the execution of the attack on the full version was discussed.
4.1 An illustrative attack on a scaled-down instance
An instance of the cipher using LFSRs described by the following primitive polynomials was used for the attack:
  • $L_{1}:f_{1}(x)=x^{3}+x+1,$
  • $L_{2}:f_{2}(x)=x^{3}+x+1,$
  • $L_{3}:f_{3}(x)=x^{3}+x+1,$
  • $L_{4}:f_{4}(x)=x^{3}+x+1.$
It should be emphasized that to simplify the generated problem as much as possible the registers of the scaled cipher are of equal length. For the selected cipher instance, the following keystream was used:
$z=(0,0,1,0,1,1,1,1,0,0,1,0).$ (24)
For the given sequence, 96 equations were generated. Based on the equations, a QUBO problem with 240 variables was determined. The problem was transformed to an Ising model and encrypted using the idea presented in [8]. The relevant part of the used key is presented as equation (25).
[…]
An oracle was built based on [21]. Using it, a parameterized problem was generated according to the idea in the section 3.2. From the parameterized vector of biases, the equations presented as Equation (26) are selected to recover the keystream. Out of 240 equations, 12 equations were selected for further analysis.
$\begin{cases}f_{121}:-z_{0}-1=1,\\ f_{131}:-z_{1}-1=1,\\ f_{141}:-z_{2}-1=2,\\ f_{151}:-z_{3}-1=-1,\\ f_{161}:-z_{4}-1=2,\\ f_{171}:-z_{5}-1=2,\\ f_{181}:-z_{6}-1=-2,\\ f_{191}:-z_{7}-1=2,\\ f_{201}:-z_{8}-1=1,\\ f_{211}:-z_{9}-1=-1,\\ f_{221}:-z_{10}-1=-2,\\ f_{231}:-z_{11}-1=1.\end{cases}$ (26)
The equation analysis is shown in Table 1. Based on it, the coefficients with changed signs were determined.
Table 1: Analysis of selected equations. Determination of part of the key bits.
$i$ Equation $z_{i}\in\{0,1\}$ $x_{i}$
121 $-z_{0}-1=1\to z_{0}=-2$ 1
131 $-z_{1}-1=1\to z_{1}=-2$ 1
141 $-z_{2}-1=2\to z_{2}=-3$ 1
151 $-z_{3}-1=-1\to z_{3}=0$ 0
161 $-z_{4}-1=2\to z_{4}=-3$ 1
171 $-z_{5}-1=2\to z_{5}=-3$ 1
181 $-z_{6}-1=-2\to z_{6}=1$ 0
191 $-z_{7}-1=2\to z_{7}=-3$ 1
201 $-z_{8}-1=1\to z_{8}=-2$ 1
211 $-z_{9}-1=-1\to z_{9}=0$ 0
221 $-z_{10}-1=-2\to z_{10}=1$ 0
231 $-z_{11}-1=1\to z_{11}=-2$ 1
Finally, based on the above analysis, the keystream was determined. The entire concealment key was recovered using the remaining constant coefficients of the matrix. The key is shown in the Figure 5.
Figure 5: Key recovered as a result of the experiment.
Missing bits, denoted as $x$ in the Figure 5, of the key do not affect the ciphertext, it is not possible to change the sign of the coefficient equal to 0. A comparison of the recovered keystream and the key confirms the correctness of the attack.
4.2 Attack on the full version of the $E_{0}$ cipher
An attack on an encrypted problem corresponding to the cryptanalysis of the complete cipher $E_{0}$ can be performed similarly. The problem has 3071 variables, and the wanted keystream has 129 bits. As in the rescaled example, the longest step is the parameterization of the matrix. The time required to generate the parameterized matrix may be estimated for about 13.5 hours. It should be noted that the script used is not optimized in any way, and the experiment only illustrates the disparity between the time of the actual phase of the attack and the time of the pre-computations. The attack phase requires solving only 129 affine equations of one variable.
As can be seen from the above time disparity, once the longer pre-calculations are performed once, the time required to execute the attack is practically imperceptible.
5 Summary and future work
This paper shows a practical attack on a proposed scheme to ensure the privacy of problems sent to quantum computing clouds. In addition, a practical attack on a small version of the cipher was demonstrated to show the correctness of the attack. It should be noted that the work does not address the correctness of the scheme under attack. The proposed attack was prepared with assumptions in the original work and additional ones resulting from the functioning of available services.
Further work should look for a secure way to secure cloud computing based on the Ising model. As an alternative to using an untrusted provider, methods can be developed that allow local, using private infrastructure, to solve the problems discussed.